OpenStudy (anonymous):

If given f(x) = –4x – 2 and g(x) = 2x, what is (f · g)(x)?

OpenStudy (anonymous):

\[f(x)=-4x-2\]\[g(x)=2x\]\[f(g(x)=-4(2x)-2=-8x-2\]

OpenStudy (anonymous):

what's the answer?

OpenStudy (anonymous):

Exactly what I showed you\[f(g(x))=-8x-2\]